# Neutrino flux sensitivity to the next galactic core-collapse supernova in COSINUS

### v1

@@ Abstract
While neutrinos are often treated as a background for many dark matter experiments, these particles offer a new avenue for physics: the detection of core-collapse supernovae. Supernovae are extremely energetic, violent and complex events that mark the death of massive stars. During their collapse stars emit a large number of neutrinos in a short burst. These neutrinos carry 99% of the emitted energy which makes their detection fundamental in understanding supernovae. This paper illustrates how COSINUS (Cryogenic Observatory for SIgnatures seen in Next-generation Underground Searches), a sodium iodide (NaI) based dark matter search, will be sensitive to the next galactic core-collapse supernova. The experiment is composed of two separate detectors which will be sensitive to far away and nearby supernovae. The inner core of the experiment will consist of NaI crystals operating as scintillating calorimeters. These crystals will mainly be sensitive to the Coherent Elastic Scattering of Neutrinos (CENS) against Na and I nuclei. The low mass of the cryogenic detectors enables the experiment a sensitivity to close supernovae below 1 kpc without pileup. The crystals will see up to hundreds of CENS events from a supernova happening at 200 pc. They reside at the center of a large cylindrical 230 T water tank, instrumented with 30 photomultiplier tubes. This tank acts simultaneously as a passive and active shield able to detect the Cherenkov radiation induced by impinging charged particles from ambient and cosmogenic radioactivity. A supernova near the Milky Way Center (10 kpc) will be easily detected inducing 60 measurable events, and the water tank will have a 3 sensitivity to supernovae up to 22 kpc, seeing 10 events. This paper shows how, even without dedicated optimization, modern dark matter experiments will also be able to play their part in the multi-messenger effort to detect the next galactic core-collapse supernova.
@@ 1 Introduction
Triggered by its own gravitational collapse, the death of a massive star is marked by the emission of a violent shock-wave (core-collapse supernova) and the birth of new compact objects, such as neutron stars or black holes. These supernovae are some of the most intense sources of neutrinos in our universe, as over a timescale of a few seconds, 99 of the stellar gravitational binding energy is converted into neutrinos with energies on the order of a few tens of MeV [1]. SN1987A was the first supernova from which neutrinos were observed, but only about two dozen events where detected from this astrophysical phenomenon [2]. The rate of galactic core-collapse supernovae has been predicted to be 1.63 0.46 per century, but this rate is not uniform throughout the galaxy [3]. One of the stars closest to earth that is likely to undergo a core-collapse supernova is Betelgeuse which is located at 200 pc [4].
@@ 2 Experimental Setup
A water tank is used as a passive and active shield for the experiment. During operation, the dilution fridge is lowered into a drywell so that the detector modules are surrounded by water without the outside of the cryostat getting wet. This drywell is 3.66 m deep and has a radius of 0.33 m. The water tank itself is 7 m tall with a diameter of 7 m and is filled with ultra-pure water. The active veto is accomplished with 30 photomultiplier tubes (PMTs) arranged inside the tank. The PMTs are 8-inch diameter R5912-30 from Hamamatsu [38]. Their arrangement can be seen on the floor and walls of the water tank in Fig. 1. To account for the physical space of the PMTs and to reduce triggers from ambient gamma rays, a 30 cm dead layer was introduced near the walls of the water tank. Tyvek, a highly reflecting material, is hung with 30 cm gap from the wall providing the dead layer and an improving light collection. The type of Tyvek used is 1082D, with two layers of Tyvek sandwiching a polyethylene layer. The reflectivity of this foil has been measured to be at least 95% [39]. The properties needed for the foil and the positioning of the PMTs were investigated using Monte Carlo simulations. The tank was optimized to be as efficient as possible as a muon veto [36].
@@ 3.2.1 Neutrino Interaction Channels in NaI cryogenic calorimeters
where is the number of target nuclei and is the neutrino flux summed over all flavors defined by Eqn. 3.1. The minimum energy required for a target nucleus to recoil is
where is the recoil energy of the nucleus, and is the target nucleus mass. The standard model differential cross-section [50, 51, 52] for neutrino scattering can be written as
where is the Fermi constant, and are the number of protons and neutrons of the target nucleus, is the weak mixing angle and is the form factor. The Klein-Nystrand [53] form factor was used for this analysis and is defined as
where , j1 is the spherical Bessel function of order one, is the range of the Yukawa potential (0.7 fm [54]) and . is the root-mean square neutron (proton) radius and is taken as 2.92 fm (2.99 fm) for 23Na [55, 56] and 4.84 fm (4.75 fm) for 127I [57, 56].
An additional channel of neutrino detection available in NaI crystals is charged current (CC) interactions with 127I, defined as
@@
The neutrino energy threshold for lepton production will be 1 MeV, and the electron and positron produced will be visible in the e-/ band. Given the large mass of the Te/Xe nuclei most of the neutrino energy becomes the lepton’s kinetic energy ( MeV). Detector efficiencies at such energies will be close to 100. Charged current interactions with or neutrinos are not possible as the energy of the supernova neutrinos will be below the required production threshold. Fig. 3 compares the different interaction cross-section of neutrinos at supernova energies in NaI. The CENS cross-sections are calculated directly from Eqn. 3.4111Where knowledge of physical parameters is required (e.g., ), the values are taken from [58] unless otherwise specified. while the charged current cross-section are taken from SNOwGLoBES [59, 60]. It is clear from Fig. 3 that within the NaI crystals the CENS channel in 127I will dominate the neutrino interactions by more than an order of magnitude. There are other channels of neutrino interactions available for the NaI crystals (neutral current (NC) with 23Na or 127I or charged current with 23Na) but either the interaction cross-section will be too small and no interactions would occur.
@@ 3.2.2 Neutrino Interaction Channels in the Water Cherenkov Detector
For most of the detectable processes222Estimation of electron scattering in the water tank from supernova neutrinos cannot be analyzed using Eqn. 3.8, and will behave similarly to CENS as described in Sec. 3.2.1. Appendix A gives the full details for this process. listed in Table 1 the number of detected supernova events can be given by [63]
where is the number of targets for the interaction , is the supernova neutrino flux defined in Sec. 3.1, is the interaction cross-section, is the detection efficiency and is the threshold energy defined in Table 1. For the energy of supernova neutrinos the cross-sections of the interactions of Table 1 are shown in Fig. 4, with the data taken from SNOwGLoBES [59, 60, 64]. It is clear that for the average expected energy of supernova neutrinos (10 - 20 MeV) inverse beta decay will be the primary channel of neutrino interaction. However, detector efficiency can increase or decrease the prominence of certain interaction channels, which is discussed in the following section.
@@ 3.2.3 Water Cherenkov Detector Efficiency
Electrons, positrons and -rays are the detectable products of neutrino interactions within the water tank. As described in Sec. 2.1, the COSINUS water tank was modeled in Geant4 to evaluate the detection efficiency of these particles over a typical supernova energy range (1 - 100 MeV). The simulation was performed by selecting the energy and particle type and uniformly distributing it inside the water tank volume. The particle momentum was isotropic, and the amount of light that arrived in the PMTs was recorded. Different trigger conditions (i.e. PMT threshold, PMTs in coincidence, etc.) could then be evaluated to determine if the particle was detected. The efficiency is determined as the number of detected particles over the total number of particles simulated. This process was then repeated over the above mentioned energy range. Fig. 5 shows the energy dependent efficiency for positrons, electrons and -rays. For this analysis a single photo-electron threshold was used, as this will be the operating threshold of the COSINUS PMTs [36], and the quantum efficiency of the PMTs was selected to be 30, corresponding to the specifications of the manufacturer at the Cherenkov wavelength. Fig. 5 also shows how the efficiency varies depending on the number of PMTs that are required to trigger for an event to be classified as detected. Based on concerns of high-background triggers the COSINUS experiment will require 4 - 6 PMTs to trigger during run-time conditions. The efficiency for detecting positrons and electrons was found to be practically identical, while the -ray efficiency was about 10 less. Specifically, the efficiency for detecting the 2.2 MeV -ray from a neutron capture was found to be 0.071, making this channel unviable for this experiment.
3.3 Neutrino Oscillation
Neutral current scattering processes such as CENS or electron scattering are flavor independent and neutrino oscillation effects will be negligible. However, charged current processes like inverse beta decay, are sensitive to only specific flavors of neutrinos and oscillation effects must be taken into account. Incoming and may change to and , respectively, (or vice-versa) during their propagation to Earth. Following the process described by [65, 66, 67], which considers vacuum oscillations, and the Mikheyev-Smirnov-Wolfenstein effect in the stellar envelope but not Earth-matter effects, the and flux reaching earth can be described by
and
where is the initial neutrino flux defined by Eqn. 3.1 and and are the survival probabilities. The survival probabilities will be dependent on the neutrino mass hierarchy and are given as
and
@@ 4.1 Detection in NaI Cryogenic Calorimeters
Assuming an average duration for a supernova of 10 s, that CENS interactions are uniformly distributed in the number of modules for each target mass, and that the time window around a remoTES cryogenic event is 100 ms [37] the probability of a signal event arriving within the same window as another event can be determined. In Fig. 7 the pileup events are removed from the total expected events, resulting in the bending of the CENS event distribution seen at very near distances. Ultimately, due to the modular nature of the COSINUS NaI, the effect of pileup is only noticeable at supernova distance less then 0.1 kpc.
The expected number of CC127I events was also investigated using the cross-section shown in Fig. 3 and Eqn. 3.8. Unfortunately, the number of events will not be significant at any relevant supernova distances. For example, for the COSINUS-2 phase, at the Betelgeuse distance of 0.197 kpc, only 1 event is expected to be observed. However, knowledge of the CC127I cross-section is fairly limited with the COHERENT experiment just recently publishing the first measurement of the energy-inclusive cross-section [70]. Therefore, in the event of a supernova, the e-/ band should still be analyzed in search of this interaction channel.
@@ 4.2 Detection in Water Cherenkov Detector
The sheer size of the water tank will allow for a supernova sensitivity to much larger distances than the crystals which reside inside. The number of expected events in the water tank can be determined by applying Eqn. 3.8, the cross-sections in Fig. 4, the efficiencies in Fig. 5 and the effect of neutrino oscillation (discussed in Sec. 3.3) to the available interactions in Table 1. The exception is the electron recoil interactions, which are detailed in Appendix A. For a supernova at 10 kpc, Table 2 gives the number of expected events per interaction type, for normal and inverse ordering scenarios and for a PMT triggering requirement of 4 and 6-fold. Oxygen NC interactions are given as upper limits due to the small number of expected events. The error given is purely statistical, as it was found that systematic errors of the efficiency and required measured parameters were negligible. As expected, it is clear from this table that inverse beta decay events will be the dominant contribution to the total number of neutrino events.
@@ 5 Conclusion
The COSINUS dark matter experiment has been evaluated for its sensitivity to the next galactic core-collapse supernova. Through modular, cryogenic, NaI calorimeters and a large water Cherenkov detector the COSINUS experiment can probe supernovae both near and far in the Milky Way. With CENS the NaI crystals will be sensitive to all flavors of neutrinos from supernovae up to 1 kpc. Their modular nature minimizes the effects of pileup and the ability for event-by-event particle discrimination creates a background free environment during a supernova explosion. Supernova neutrino interactions in the water tank will be dominated by inverse beta decay events and for the assumptions considered in this study the COSINUS water tank could be sensitive to a supernova between 15 and 22 kpc away. Future work will involve integrating the COSINUS experiment into the Supernova Early Warning System [72] and evaluating the directional sensitivity of the water Cherenkov detector. Moreover, as discussed in [47], COSINUS can be able to identify the time of black-hole formation in case of Failed CCSNe with an uncertainty of a few ms, providing a useful alert for the nearby gravitational wave interferometer Virgo to look for the GW counterpart. Finally, the COSINUS experiment, with one large homogeneous detector containing modular cryogenic calorimeters in the center, is distinctly sensitive to almost all of the possible neutrino interactions in matter (CENS, CC, NC, and electron scattering). Although COSINUS was not optimized for the detection of supernova neutrinos, the variety of interaction channels it is sensitive to and the possibility of coincidence analysis could represent a design philosopy to follow for future dedicated supernova neutrino experiments.
Appendix A Electron-Neutrino Scattering
As discussed in Sec. 3.2.2 neutrinos can scatter off the electrons present in the water tank. Traditionally, this interaction can provide some directionality with regard to the location of the supernova. This process is sensitive to all flavors of neutrinos and the differential recoil rate is described as
where is the number of target electrons and is the neutrino flux defined by Eqn. 3.1. Based of conservation of momentum, the minimum neutrino energy to generate a electron recoil of energy can be written as
where is the mass of the electron. From [73] the differential cross-section can be written as
with the coefficients and being different for each flavor of neutrino, see Table 4.
sin | sin | - sin | sin | |
sin | sin | sin | - sin |
For a supernova at 10 kpc the differential recoil spectrum is shown in Fig. 9. Both the normal and inverse ordering scenario are shown, as well the efficiency from Fig. 5 is folded in. By integrating the spectra and multiplying by the water tank mass the total number of electron scattering events can be calculated. Table 5 shows the number of events for a supernova at 10 kpc. Repeating the above process for different supernovae distances a profile of the total number of electron scattering events can be built, shown in Fig. 10. As shown, electron scattering will contribute only a small percentage of the total number of neutrino interactions and will only be relevent for very close supernovae.
Flux | Ordering | PMT Threshold | Total | ||||
1987A | NO | 4 | 1.64 | 0.17 | 0.31 | 0.34 | 2.46 |
1987A | IO | 4 | 1.30 | 0.30 | 0.36 | 0.17 | 2.13 |
1987A | NO | 6 | 0.92 | 0.075 | 0.16 | 0.18 | 1.34 |
1987A | IO | 6 | 0.70 | 0.14 | 0.19 | 0.08 | 1.11 |
NO | 4 | 1.45 | 0.20 | 0.32 | 0.35 | 2.32 | |
IO | 4 | 1.24 | 0.30 | 0.35 | 0.29 | 2.18 | |
NO | 6 | 0.75 | 0.09 | 0.16 | 0.18 | 1.18 | |
IO | 6 | 0.63 | 0.14 | 0.18 | 0.15 | 1.10 |

### v5

@@ Abstract
While neutrinos are often treated as a background for many dark matter experiments, these particles offer a new avenue for physics: the detection of core-collapse supernovae. Supernovae are extremely energetic, violent and complex events that mark the death of massive stars. During their collapse stars emit a large number of neutrinos in a short burst. These neutrinos carry 99% of the emitted energy which makes their detection fundamental in understanding supernovae. This paper illustrates how COSINUS (Cryogenic Observatory for SIgnatures seen in Next-generation Underground Searches), a sodium iodide (NaI) based dark matter search, will be sensitive to the next galactic core-collapse supernova. The experiment is composed of two separate detectors which will respond to far away and nearby supernovae. The inner core of the experiment will consist of NaI crystals operating as scintillating calorimeters. These crystals will mainly be sensitive to the Coherent Elastic Neutrino-Nucleus Scattering (CENS) against Na and I nuclei. The low mass of the cryogenic detectors enables the experiment to identify close supernovae within 1 kpc without pileup. The crystals will see up to hundreds of CENS events from a supernova happening at 200 pc. They reside at the center of a large cylindrical 230 T water tank, instrumented with 30 photomultiplier tubes. This tank acts simultaneously as a passive and active shield able to detect the Cherenkov radiation induced by impinging charged particles from ambient and cosmogenic radioactivity. A supernova near the Milky Way Center (10 kpc) will be easily detected inducing 60 measurable events, and the water tank will have a 3 sensitivity to supernovae up to 22 kpc, seeing 10 events. This paper shows how, even without dedicated optimization, modern dark matter experiments will also be able to play their part in the multi-messenger effort to detect the next galactic core-collapse supernova.
@@ 1 Introduction
Triggered by its own gravitational collapse, the death of a massive star is marked by the emission of a violent shock-wave (core-collapse supernova) and the birth of new compact objects, such as neutron stars or black holes. These supernovae are some of the most intense sources of neutrinos in our universe, as over a timescale of a few seconds, 99 of the stellar gravitational binding energy is converted into neutrinos with energies on the order of a few tens of MeV [1]. SN1987A was the first supernova from which neutrinos were observed, but only about two dozen events were detected from this astrophysical phenomenon [2]. The rate of galactic core-collapse supernovae has been predicted to be 1.63 0.46 per century, but this rate is not uniform throughout the galaxy [3]. One of the stars closest to earth that is likely to undergo a core-collapse supernova is Betelgeuse, which is located at 200 pc [4].
@@ 2 Experimental Setup
A water tank is used as a passive and active shield for the experiment. During operation, the dilution fridge is lowered into a drywell so that the detector modules are surrounded by water without the outside of the cryostat getting wet. This drywell is 3.66 m deep and has a radius of 0.33 m. The water tank itself is 7 m tall with a diameter of 7 m and is filled with ultra-pure water. The active veto is accomplished with 30 photomultiplier tubes (PMTs) arranged inside the tank. The PMTs are 8-inch diameter R5912-30 from Hamamatsu [38]. Their arrangement can be seen on the floor and walls of the water tank in Fig. 1. To account for the physical space of the PMTs and to reduce triggers from ambient gamma rays, a 30 cm dead layer was introduced near the walls of the water tank. Tyvek, a highly reflecting material, is hung with 30 cm gap from the wall providing the dead layer and an improving light collection. The type of Tyvek used is 1082D, with two layers of Tyvek sandwiching a polyethylene layer. The reflectivity of this foil has been measured to be at least 95% [39]. The properties needed for the foil and the positioning of the PMTs were investigated using Monte Carlo simulations. The tank was optimized to be as efficient as possible as a muon veto, achieving muon veto efficiency of 99.63 0.16 and 44.4 5.6 for showers of secondary particles (5 of all events). Giving a total efficiency of 97.0 0.3. [36].
@@ 3.2.1 Neutrino Interaction Channels in NaI cryogenic calorimeters
where is the number of target nuclei and is the neutrino flux summed over all flavors defined by Eqn. 3.1. The minimum energy required for a target nucleus to recoil is , where is the recoil energy of the nucleus, and is the target nucleus mass. The standard model differential cross-section [50, 51, 52] for neutrino scattering requires the nuclear form factor to account for incoherency at higher neutrino energies. The Klein-Nystrand [53] form factor is used for this analysis,
where is the transferred momentum, the range of the Yukawa potential is set as = 0.7 fm [54] and where is the root-mean square neutron (proton) radius and is taken as 2.92 fm (2.99 fm) for 23Na [55, 56] and 4.84 fm (4.75 fm) for 127I [57, 56].
An additional channel of neutrino detection available in NaI crystals is charged current (CC) [58] interactions with 127I, defined as
@@
The neutrino energy threshold for lepton production will be 1 MeV, and the electron and positron produced will be visible in the e-/ band. Given the large mass of the Te/Xe nuclei most of the neutrino energy becomes the lepton’s kinetic energy ( MeV). Detector efficiencies at such energies will be close to 100. Charged current interactions with or neutrinos are not possible as the energy of the supernova neutrinos will be below the required production threshold. Fig. 3 compares the different interaction cross-section of neutrinos at supernova energies in NaI. The CENS cross-sections are calculated directly from [50, 51, 52]111Where knowledge of physical parameters is required (e.g., ), the values are taken from [59] unless otherwise specified. while the charged current cross-section are taken from SNOwGLoBES [60, 61]. It is clear from Fig. 3 that within the NaI crystals the CENS channel in 127I will dominate the neutrino interactions by more than an order of magnitude. Other channels of neutrino interactions available for the NaI crystals (neutral current (NC) with 23Na or 127I or charged current with 23Na) were not considered for this study.
@@ 3.2.2 Neutrino Interaction Channels in the Water Cherenkov Detector
For most of the detectable processes listed in Table 1 the number of detected supernova events can be given by [64]
where is the number of targets for the interaction , is the supernova neutrino flux defined in Sec. 3.1, is the interaction cross-section, is the detection efficiency and is the threshold energy defined in Table 1. For the energy of supernova neutrinos the cross-sections of the interactions of Table 1 are shown in Fig. 4, with the data taken from SNOwGLoBES [60, 61, 65]. To estimate the event rate of electron-neutrino scattering using the cross section [66], an integral over both the neutrino energies and outgoing electron energies is required, similar to the procedure for calculating CENS event rates. Electron scattering will contribute only a small percentage of the total number of neutrino interactions and will only be relevant for very close supernovae. Accounting for the neutrino oscillations for all flavor dependent cross sections is done following the process described by [67, 68, 69], which considers vacuum oscillations and the Mikheyev-Smirnov-Wolfenstein effect in the stellar envelope but not Earth-matter effects. Of note is the dependence of the survival probabilities on the mass hierarchy. It is clear that for the average expected energy of supernova neutrinos (10 - 20 MeV) inverse beta decay will be the primary channel of neutrino interaction. However, detector efficiency can increase or decrease the prominence of certain interaction channels, which is discussed in the following section.
@@ 3.2.3 Water Cherenkov Detector Efficiency
Electrons, positrons and -rays are the detectable products of neutrino interactions within the water tank. As described in Sec. 2.1, the COSINUS water tank was modeled in Geant4 to evaluate the detection efficiency of these particles over a typical supernova energy range (1 - 100 MeV). The simulation was performed by choosing a monoenergetic energy (0.5, 1, 5, 10, 20, 30, 40, 50, 60, 80 or 100 MeV) and particle type (positron, electron or gamma) and uniformly distributing it inside the water tank volume. The particle momentum was isotropic, and the amount of light that arrived in the PMTs was recorded. Different trigger conditions (i.e. PMT threshold, PMTs in coincidence, etc.) could then be evaluated to determine if the particle was detected. The efficiency is determined as the number of detected events over the total number of events simulated. This process was then repeated over all the above mentioned energies and particle types. For energies in-between the simulated energy values the efficiency is interpolated between the simulated data points. Fig. 5 shows the energy dependent efficiency for positrons, electrons and -rays. For this analysis a single photo-electron threshold was used, as this will be the operating threshold of the COSINUS PMTs [36], and the quantum efficiency of the PMTs was selected to be 30, corresponding to the specifications of the manufacturer at the Cherenkov wavelength. Fig. 5 also shows how the efficiency varies depending on the number of PMTs that are required to trigger for an event to be classified as detected. Based on concerns of high-background triggers the COSINUS experiment will require 4 - 6 PMTs to trigger during run-time conditions and a coincidence window of at least 500 ns. The efficiency for detecting positrons and electrons was found to be practically identical, while the -ray efficiency was about 10 less. Specifically, the efficiency for detecting the 2.2 MeV -ray from a neutron capture was found to be 0.071, making this channel unviable for this experiment.
@@ 4.1 Detection in NaI Cryogenic Calorimeters
Assuming an average duration for a supernova of 10 s, that CENS interactions are uniformly distributed in the number of modules for each target mass, and that the time window around a remoTES cryogenic event is 100 ms [37] the probability of a signal event arriving within the same window as another event can be determined. In Fig. 7 the pileup events are removed from the total expected events, resulting in the bending of the CENS event distribution seen at very near distances. Ultimately, due to the modular nature of the COSINUS NaI, the effect of pileup is only noticeable at supernova distance less then 0.1 kpc, corresponding to 6 (8%), 24 (10%) and 270 (30%) events lost due to pileup for respectively 227 g, 816 g and 2808 g for a 1987A-like supernova.
The expected number of CC127I events was also investigated using the cross-section shown in Fig. 3 and Eqn. 3.6. Unfortunately, the number of events will not be significant at any relevant supernova distances. For example, for the COSINUS-2 phase, at the Betelgeuse distance of 0.197 kpc, only 2 CC127I events are expected to be observed. However, knowledge of the CC127I cross-section is fairly limited with the COHERENT experiment just recently publishing the first measurement of the energy-inclusive cross-section [72]. Therefore, in the event of a supernova, the e-/ band should still be analyzed in search of this interaction channel.
@@ 4.2 Detection in Water Cherenkov Detector
The sheer size of the water tank will allow for a supernova sensitivity to much larger distances than the crystals which reside inside. The number of expected events in the water tank can be determined by applying Eqn. 3.6, the cross-sections in Fig. 4, the efficiencies in Fig. 5 and the effect of neutrino oscillation to the available interactions in Table 1. The exception is the electron recoil interactions. For a supernova at 10 kpc, Table 2 gives the number of expected events per interaction type, for normal and inverse ordering scenarios and for a PMT triggering requirement of 4 and 6-fold. Oxygen NC interactions are given as upper limits due to the small number of expected events. The error given is purely statistical, as it was found that systematic errors of the efficiency and required measured parameters were negligible. As expected, it is clear from this table that inverse beta decay events will be the dominant contribution to the total number of neutrino events.
@@ 5 Conclusion
The COSINUS dark matter experiment has been evaluated for its sensitivity to the next galactic core-collapse supernova. Through modular, cryogenic, NaI calorimeters and a large water Cherenkov detector the COSINUS experiment can probe supernovae both near and far in the Milky Way. With CENS the NaI crystals will be sensitive to all flavors of neutrinos from supernovae up to 1 kpc. Their modular nature minimizes the effects of pileup and the ability for event-by-event particle discrimination creates a background free environment during a supernova explosion. Supernova neutrino interactions in the water tank will be dominated by inverse beta decay events and for the assumptions considered in this study the COSINUS water tank could be sensitive to a supernova between 15 and 22 kpc away. Future work will involve integrating the COSINUS experiment into the Supernova Early Warning System [74] and evaluating the directional sensitivity of the water Cherenkov detector. Moreover, as discussed in [47], COSINUS can be able to identify the time of black-hole formation in case of Failed CCSNe with an uncertainty of a few ms, providing a useful alert for the nearby gravitational wave interferometer Virgo to look for the GW counterpart. Finally, the COSINUS experiment, with one large homogeneous detector containing modular cryogenic calorimeters in the center, is distinctly sensitive to almost all of the possible neutrino interactions in matter (CENS, CC, NC, and electron scattering). Although COSINUS was not optimized for the detection of supernova neutrinos, the variety of interaction channels it is sensitive to and the possibility of coincidence analysis could represent a design philosophy to follow for future dedicated supernova neutrino experiments.